OpenStudy (anonymous):

Evaluate \[\lim_{x \rightarrow 1} \frac{cos(\frac{\pi x}{2})}{1-x}\]

OpenStudy (turingtest):

l'hospital allowed?

OpenStudy (anonymous):

Not allowed..

OpenStudy (turingtest):

rats

OpenStudy (experimentx):

substitute let y=x-1

OpenStudy (experimentx):

* y=1-x

OpenStudy (turingtest):

\[\lim_{y\to0}\frac{\cos(\frac\pi2(y+1))}{y}\]???

OpenStudy (turingtest):

changes to -sin okay I see it now :)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{cos(\frac{\pi x}{2})}{1-x}\]\[=\lim_{y \rightarrow 0} \frac{cos(\frac{\pi (1-y)}{2})}{y}\]\[=\lim_{y \rightarrow 0} \frac{cos(\frac{\pi }{2}-\frac{\pi y}{2})}{y}\]\[=\lim_{y \rightarrow 0} \frac{sin(\frac{\pi y}{2})}{y}\]\[=\frac{\pi}{2}\]?

OpenStudy (turingtest):

oh yeah +sin

OpenStudy (experimentx):

seems correct!! http://www.wolframalpha.com/input/?i= \lim_{x+\rightarrow+1}+\frac{cos%28\frac{\pi+x}{2}%29}{1-x}

OpenStudy (anonymous):

Yeah!~ Thanks!!~